OpenStudy (anonymous):

Two points on a line are given by the ordered pairs (0, -1) and (0, 6). What type of line is this? (Points : 3)        horizontal        vertical        slant        none of these

OpenStudy (anonymous):

slope of the line=(y2-y1)/(x2-x1) =(6-(-1))/(0-0)=7/0 \[=\infty \] Therefore it is a vertical line.

OpenStudy (anonymous):

i HAVE ANOTHER question Find the slope of the line that passes through the points (0, -3) and (-4, 1).

OpenStudy (anonymous):

you there

OpenStudy (anonymous):

Slope of required line=(y2-y1)/(x2-x1) ={1-(-3)}/(-4-0) =4/-4=-1